OpenStudy (anonymous):

A farmer has to transport 20 people and 32 sheep to a market. He can use either Fiats (x) which take 2 people and 1 sheep, or Rolls Royce (y) which take 2 people and 4 sheeps. He must not use more than 15 cars altogether. What are the inequalities ?

OpenStudy (dumbcow):

haha i can imagine driving around in a Rolls Royce with 4 sheep...

OpenStudy (dumbcow):

hilarious...anyway you know x+y <= 15

OpenStudy (anonymous):

i knew that one, but the two others equations please....

OpenStudy (dumbcow):

people: 2x +2y <= 20 sheep: x +4y <= 32

OpenStudy (anonymous):

ok thank you

OpenStudy (dumbcow):

yw